# Harnessing The Collective Wisdom: Fusion Learning Using Decision Sequences From Diverse Sources

### v2

@@ 3 IRT: more powerful evidence aggregation via multiplication
While the IRT procedure guarantees valid FDR control, the choice of $\alpha$ bears important consideration as far as its power is concerned. For instance, with a relatively smaller value of $\alpha$, IRT may fail to recover discoveries identified by studies with a smaller weight $w_{j}$. In fact, when inferences are pooled with the goal of achieving higher reliability then often $\alpha<\min_{j\in[d]}\alpha_{j}$, and in such settings IRT may exhibit no power. In Section 3.1 we take a simple example to discuss the impact that $\alpha$ has on the power of IRT. Thereafter, in Section 3.2 we present IRT, which relies on an alternative evidence aggregation scheme and is more powerful than IRT when inferences are synthesized for higher reliability.
3.1 Impact of $\alpha$ on the power of IRT
@@
Suppose there are $d=2$ studies, each testing the same set of $m=5$ null hypotheses at levels $\alpha_{1}$ and $\alpha_{2}$, respectively. Consider a simple setting where the $d$ studies reject only the $i$-th null hypothesis. So $\delta_{ij}=1$ and $\|\bm{\delta}_{j}\|_{0}=1$ for all $j\in[d]$. Suppose IRT is used to pool inferences from these studies. The gray shaded region in the left panel of Figure 2 depicts the overall FDR level $\alpha$ required for the $e-$BH procedure to reject $H_{0i}$ when $\alpha_{1}$ varies over $0.001$ to $0.1$ and $\alpha_{2}$ is fixed at $0.05$. Here the red dotted line represents $\min(\alpha_{1},\alpha_{2})$. Most notably, this plot reveals that one must have $\alpha>\min(\alpha_{1},\alpha_{2})$ to reject $H_{0i}$ unless $\alpha_{1}=\alpha_{2}=0.05$, in which case $\alpha$ must at least be $0.05$ for the $e-$BH procedure to reject $H_{0i}$. The right panel considers the same setting but with $d=3,\alpha_{3}=0.03$ and paints a similar picture.
@@
Figure 2: Pooling inferences using IRT. Left: Here $d=2$ studies are testing the same set of $m=5$ null hypotheses at levels $\alpha_{1}$ and $\alpha_{2}$, respectively. Both studies reject only the $i^{\tt th}$ null hypothesis. The gray shaded region depicts the overall FDR level $\alpha$ required for the $e-$BH procedure to reject $H_{0i}$ when $\alpha_{1}$ varies over $0.001$ to $0.1$ and $\alpha_{2}$ is fixed at $0.05$. Here the red dotted line represents $\min(\alpha_{1},\alpha_{2})$. Right: Same setting with $d=3$ and $\alpha_{3}=0.03$.
The calculations for Figure 2 readily follow from the IRT procedure. For instance, in the case of the left panel, IRT rejects $H_{0i}$ at FDR level $\alpha$ if $\alpha\geq 2\alpha_{1}\alpha_{2}/(\alpha_{1}+\alpha_{2})$. Furthermore, noting that the quantity on the right side of the inequality is the harmonic mean of $\{\alpha_{1},\alpha_{2}\}$, we have $2\alpha_{1}\alpha_{2}/(\alpha_{1}+\alpha_{2})\geq\min(\alpha_{1},\alpha_{2})$ where equality holds when $\alpha_{1}=\alpha_{2}$. However, the results from the left panel are surprising, since intuitively the decisions of the second study should enhance the evidence against $H_{0i}$. How can it be that the evidence from the first study is sufficient to reject $H_{0i}$, but the combined evidence from the two studies is not sufficient to reject $H_{0i}$ unless $\alpha>\min(\alpha_{1},\alpha_{2})$? The reason behind this seemingly counterintuitive situation is that IRT makes no assumption on the dependence between the two studies. For instance, the two studies could be employing identical data and testing methods to reach the same conclusion. Then the collective evidence should not necessarily be interpreted as stronger than the individual evidence from each study, given that the latter study does not introduce any fresh information. For the evidence to truly accumulate, we often presume a certain level of difference in the data or testing procedures applied in these studies. However, such an assumption may be unwarranted when only binary decisions and specified FDR levels are available. Absent any dependence assumptions, the most logical approach would be to remain conservative in our interpretations. The IRT aggregation step described in Section 2.2 follows this principle and aggregates the evidence indices via their arithmetic mean. In fact, it is demonstrated in Vovk and Wang (2021) that without further assumptions on the dependence between $e-$values, the arithmetic mean essentially dominates all other $e-$value aggregation methods.
3.2 Alternative evidence aggregation scheme
@@
Note that the notion of partial exchangeablilty on the set of nulls is commonly used in the conformal inference literature (Bates et al., 2023; Liang et al., 2024) while symmetric decision rules arise naturally in conventional settings where all hypotheses undergo simultaneous testing without the inclusion of auxiliary side information. Lemma 1 guarantees that if for each study the summary statistics are partially exchangeable and the testing procedure is symmetric then a scaled version of the evidence indices in Equation (1) are actually bonafide $e-$values.
@@
Proof of Lemma 1..
@@
Bonafide $e-$values are substantially more flexible than generalized $e-$values. For instance, $e-$values facilitate reliable inferences for individual hypotheses while generalized $e-$values are limited to simultaneous inference across a set of null hypotheses. Furthermore, the only aggregation scheme for generalized $e-$values discussed in the literature is (weighted) arithmetic mean (Ren and Barber, 2024; Li and Zhang, 2023). In contrast, under independence bonafide $e-$values admit aggregation via multiplication, which allows evidence to “accumulate”.
Let $\mathcal{N}_{i}=\{j:\mathbb{I}(i\in\mathcal{M}_{j})=1\}$ denote the set of studies that test hypothesis $H_{0i}$ with $|\mathcal{N}_{i}|=n_{i}$. Given a pre-determined $k\in\{1,\ldots,n_{i}\}$, denote $\mathcal{S}_{ki}$ as any $k$ element subset of $\mathcal{N}_{i}$, $i\in\mathcal{M}$. Let $\pi\in(0,1)$ be a lower bound of $\min_{j\in[d]}\{|\mathcal{H}_{0j}|/m_{j}\}$ and define $A_{i,\mathcal{S}_{ki}}=\prod_{j\in\mathcal{S}_{ki}}e_{ij}$. The next theorem shows that under certain conditions $e_{i}=\pi^{k}A_{i,\mathcal{S}_{ki}}$ is an $e-$value.
Theorem 3.
Suppose (i) the conditions in Lemma 1 hold and (ii) the summary statistics for the $i$-th testing problem, $\{X_{ij}\}_{j\in\mathcal{N}_{i}}$ are independent conditional on $\theta_{i}$ for all $i$. Let $e_{i}=\pi^{k}A_{i,\mathcal{S}_{ki}}$. Then $\mathbb{E}[e_{i}]\leq 1$ for $i\in\mathcal{H}_{0}$.
Proof of Theorem 3..
Note that $\mathbb{E}\left[A_{i,\mathcal{S}_{ki}}\right]=\prod_{j\in\mathcal{S}_{ki}}\mathbb{E}[e_{ij}]\leq\prod_{j\in\mathcal{S}_{ki}}\dfrac{m_{j}}{|\mathcal{H}_{0j}|}$ and thus $\mathbb{E}[\pi^{k}A_{i,\mathcal{S}_{ki}}]\leq 1$. ∎
Theorem 3 facilitates evidence accumulation via a product rule as we can multiply the $\{e_{ij}\}_{j\in\mathcal{\mathcal{missing}}S_{ki}}$’s for aggregation. However, simply multiplying them may not be ideal since if just one study in $\mathcal{S}_{ki}$ fails to reject $H_{0i}$ the product will be 00. To partially over come this difficulty, we propose to use $e_{i}^{\tt agg*}$ as defined below for evidence aggregation.
$e_{i}^{\tt agg*}=\frac{1}{n_{i}}\sum_{k=1}^{n_{i}}\binom{n_{i}}{k}^{-1}\sum_{\mathcal{S}_{ki}\in\mathcal{B}_{ki}}\pi^{k}A_{i,\mathcal{S}_{ki}},$ (7)
where $\mathcal{B}_{ki}$ is the set of all $k$ element subsets of $\mathcal{N}_{i}$. The idea is to try all possible $\pi^{k}A_{i,\mathcal{S}_{ki}}$ and then take average. Since $\pi^{k}A_{i,\mathcal{S}_{ki}}$ are $e-$values, their average is also an $e-$value. We denote the procedure that applies the $e-$BH procedure on $(e_{1}^{\tt agg*},\ldots,e_{m}^{\tt agg*})$ as IRT. In the real data analysis and numerical experiments of sections 5 and 6 we take $\pi_{j}=0.5$ for all $j\in[d]$, as is often assumed in the literature (Jin and Cai, 2007).
We now return to the example considered in Section 3.1 but instead use IRT to pool the inferences from the two studies. The green and the gray shaded regions in the left panel of Figure 3 depicts the overall FDR level $\alpha$ required for the $e-$BH procedure to reject $H_{0i}$ as a function of $\alpha_{1}$ and with $\alpha_{2}=0.05$. Clearly, IRT is able to reject $H_{0i}$ even when $\alpha<\min(\alpha_{1},\alpha_{2})$. This represents a stark contrast to IRT which requires $\alpha$ to stay in the gray region to reject $H_{0i}$. The reason for this distinction is related to the fact that, courtesy Theorem 3, the aggregation scheme underlying IRT is more powerful than the arithmetic mean. Note that the calculations for the left panel directly follow from the IRT procedure which, in this example, rejects $H_{0i}$ at FDR level $\alpha$ if $\alpha\geq 8\alpha_{1}\alpha_{2}/(\alpha_{1}+\alpha_{2}+m)$. While this example relies on a relatively simple setting, the numerical experiments in Section 6 confirm the broader conclusion that, in general, IRT is not powerful when $\alpha<\min_{j\in[d]}\alpha_{j}$ while, under the conditions of Theorem 3, IRT is powerful in this setting.
Remark 1.
The preceding discussion naturally begs the question regarding the practical relevance of IRT while pooling inferences for higher reliability. Here, we address this concern.
Motivated by the setting in the left panels of figures 2 and 3, suppose there are $d=4$ studies, each testing the same set of $m=5$ null hypotheses. The first two studies are testing at level $\alpha_{1}$ while the remaining two at $\alpha_{2}$. Consider a setting where the first two studies are using exactly the same data and both reject only the $i$-th null hypothesis at FDR level $\alpha_{1}$. Other two studies are independent such that Theorem 3 holds, but both continue to reject the same $i$-th null hypothesis at FDR level $\alpha_{2}$. So $\delta_{ij}=1$ and $\|\bm{\delta}_{j}\|_{0}=1$ for all $j\in[d]$. To pool inferences across these $d$ studies, IRT and IRT can be used in conjunction. Specifically, the aggregated evidence indices from the first two studies are constructed using IRT, denoted $e_{i}^{\tt agg}$ (Equation (3)) and IRT is employed to aggregate the evidence indices from the remaining two studies, denoted $e_{i}^{\tt agg,*}$ (Equation (7)). Denote $e_{i}^{\tt agg,H}=(1/2)(e_{i}^{\tt agg}+e_{i}^{\tt agg,*})$ as the hybrid aggregated evidence indices. Note that $\{e_{i}^{\tt agg,H}\}_{i=1}^{m}$ are a set of generalized $e-$values under $\mathcal{H}_{0}$. When $\{e_{i}^{\tt agg,H}\}_{i=1}^{m}$ are used as inputs to the $e-$BH procedure, the gray and the green shaded regions in the right panel of Figure 3 depict the overall FDR level $\alpha$ required for $e-$BH to reject $H_{0i}$ as a function of $\alpha_{1}$ and with $\alpha_{2}=0.05$. Clearly, this hybrid scheme is able to reject $H_{0i}$ even when $\alpha<\min(\alpha_{1},\alpha_{2})$. In contrast, the gray shaded region reveals that IRT has no power unless $\alpha>\min(\alpha_{1},\alpha_{2})$.
The above example represents a practical setting where often additional information regarding data-sharing or the use of auxiliary side information for multiple testing is available for the $d$ studies. For instance, suppose prior knowledge dictates that a set of $d_{1}\subset[d]$ studies share data amongst themselves, while the conditions of Theorem 3 hold for the remaining set of $d_{2}=[d]\setminus d_{1}$ studies. Denote the aggregated evidence indices across the $d_{1}$ studies derived from IRT as $\{e_{i}^{\tt agg,d_{1}}\}_{i\in\mathcal{M}}$ and those derived from IRT across the $d_{2}$ studies as $\{e_{i}^{\tt agg*,d_{2}}\}_{i\in\mathcal{M}}$. Then, the hybrid aggregated evidence indices $e_{i}^{\tt agg,H}=(1/d)(|d_{1}|e_{i}^{\tt agg,d_{1}}+|d_{2}|e_{i}^{\tt agg*,d_{2}})$ are a set of generalized $e-$values under $\mathcal{H}_{0}$ and the $e-$BH procedure guarantees FDR control at level $\alpha$ when $\{e_{i}^{\tt agg,H}\}_{i=1}^{m}$ are used as inputs. We call this procedure IRT H and evaluate its empirical performance in Section 6.
4 Discussion
We present additional insights about the IRT framework. In particular, we (i) show that the evidence indices in Equation (1) are connected to some existing aggregation and derandomization procedures (Section 4.1), (ii) prove that IRT guarantees asymptotic FDR control if some studies control their FDR asymptotically (Section 4.2), and (iii) develop a procedure, along the lines of IRT, for pooling inferences when the study-specific $p-$values are available (Section 4.3).
4.1 Connections to existing aggregation and derandomization procedures

### v3

@@ Appendix A Impact of $\alpha$ on the power of IRT
When all $d$ studies report the triplets $\{\mathcal{D}_{j}:j\in[d]\}$, the choice of $\alpha$ bears important consideration as far as the power of IRT is concerned, where we refer to IRT as the procedure that applies the e-BH method directly to the aggregated e-values, $\boldsymbol{e}^{\text{agg}}$. For instance, with a relatively smaller value of $\alpha$, IRT may fail to recover discoveries identified by studies with a smaller weight $w_{j}$. In fact, when inferences are pooled with the goal of achieving higher reliability then often $\alpha<\min_{j\in[d]}\alpha_{j}$, and in such settings IRT may exhibit no power. In this section we take a simple example to discuss the impact that $\alpha$ has on the power of IRT. Thereafter, in Section B we present IRT, which relies on an alternative evidence aggregation scheme and is more powerful than IRT when inferences are synthesized for higher reliability.
@@
Suppose there are $d=2$ studies, each testing the same set of $m=5$ null hypotheses at levels $\alpha_{1}$ and $\alpha_{2}$, respectively. Consider a simple setting where the $d$ studies reject only the $i$-th null hypothesis. So $\delta_{ij}=1$ and $\|\bm{\delta}_{j}\|_{0}=1$ for all $j\in[d]$. Suppose IRT is used to pool inferences from these studies. The gray shaded region in the left panel of Figure 3 depicts the overall FDR level $\alpha$ required for the e-BH procedure to reject $H_{0i}$ when $\alpha_{1}$ varies over $0.001$ to $0.1$ and $\alpha_{2}$ is fixed at $0.05$. Here the red dotted line represents $\min(\alpha_{1},\alpha_{2})$. Most notably, this plot reveals that one must have $\alpha>\min(\alpha_{1},\alpha_{2})$ to reject $H_{0i}$ unless $\alpha_{1}=\alpha_{2}=0.05$, in which case $\alpha$ must at least be $0.05$ for the e-BH procedure to reject $H_{0i}$. The right panel considers the same setting but with $d=3,~\alpha_{2}=0.05,~\alpha_{3}=0.03$ and paints a similar picture.
@@
Figure 3: Pooling inferences using IRT. Left: Here $d=2$ studies are testing the same set of $m=5$ null hypotheses at levels $\alpha_{1}$ and $\alpha_{2}$, respectively. Both studies reject only the $i^{\tt th}$ null hypothesis. The gray shaded region depicts the overall FDR level $\alpha$ required for the e-BH procedure to reject $H_{0i}$ when $\alpha_{1}$ varies over $0.001$ to $0.1$ and $\alpha_{2}$ is fixed at $0.05$. Here the red dotted line represents $\min(\alpha_{1},\alpha_{2})$. Right: Same setting with $d=3,~\alpha_{2}=0.05$ and $\alpha_{3}=0.03$.
The calculations for Figure 3 readily follow from the IRT procedure. For instance, in the case of the left panel, IRT rejects $H_{0i}$ at FDR level $\alpha$ if $\alpha\geq 2\alpha_{1}\alpha_{2}/(\alpha_{1}+\alpha_{2})$. We have $2\alpha_{1}\alpha_{2}/(\alpha_{1}+\alpha_{2})\geq\min(\alpha_{1},\alpha_{2})$ where equality holds when $\alpha_{1}=\alpha_{2}$. The results from the left panel may seem counter-intuitive, as the decisions of the second study do not necessarily enhance the evidence against $H_{0i}$. However, this in fact reflects a fundamental constraint inherent to the nature of the problem. If the two studies use identical data and methods, the collective evidence may not be stronger than the individual evidence, since the latter study does not provide fresh information. No fusion learning method, including IRT, could justifiably claim a rejection at a lower FDR level than $\min\{\alpha_{1},\alpha_{2}\}$. The IRT framework is designed with the explicit goal of guaranteeing FDR control under minimal assumptions. This principle requires the framework to be conservative enough to remain valid even in worst-case scenarios of inter-study dependence.
Appendix B IRT: more powerful evidence aggregation via multiplication
@@
The notion of partial exchangeablilty on the set of nulls is commonly used in the conformal inference literature (Bates et al., 2023; Liang et al., 2024) while symmetric decision rules arise naturally in conventional settings where all hypotheses undergo simultaneous testing without the inclusion of auxiliary side information. Lemma 1 guarantees that if for each study the summary statistics are partially exchangeable and the testing procedure is symmetric then a scaled version of the evidence indices in Equation (3) are actually bonafide $e$-values.
@@
$e$-values are substantially more flexible than compound $e$-values. For instance, $e$-values facilitate reliable inferences for individual hypotheses while compound $e$-values are limited to simultaneous inference across a set of null hypotheses. Furthermore, the only aggregation scheme for compound $e$-values discussed in the literature is (weighted) arithmetic mean (Ren and Barber, 2024; Li and Zhang, 2023). In contrast, under independence bonafide $e$-values admit aggregation via multiplication, which allows evidence to “accumulate”. Let $\mathcal{N}_{i}=\{j:\mathbb{I}(i\in\mathcal{M}_{j})=1\}$ denote the set of studies that test hypothesis $H_{0i}$ with $|\mathcal{N}_{i}|=n_{i}$. Given a pre-determined $k\in\{1,\ldots,n_{i}\}$, denote $\mathcal{S}_{ki}$ as any $k$ element subset of $\mathcal{N}_{i}$, $i\in\mathcal{M}$. Define $e_{i,\mathcal{S}_{ki}}=\prod_{j\in\mathcal{S}_{ki}}\pi_{j}e_{ij}$. The next theorem shows that under certain conditions $e_{i,\mathcal{S}_{ki}}$ is an $e$-value.
Theorem 3.
Suppose (i) the conditions in Lemma 1 hold and (ii) the summary statistics for the $i$-th testing problem, $\{X_{ij}\}_{j\in\mathcal{N}_{i}}$ are independent conditional on $\theta_{i}$ for all $i$. Then $\mathbb{E}(e_{i,\mathcal{S}_{ki}})\leq 1$ for $i\in\mathcal{H}_{0}$.
Theorem 3 facilitates evidence accumulation via a product rule as we can multiply the $\{e_{ij}\}_{j\in\mathcal{\mathcal{missing}}S_{ki}}$’s for aggregation. However, simply multiplying them may not be ideal since if just one study in $\mathcal{S}_{ki}$ fails to reject $H_{0i}$ the product will be 0. To partially over come this difficulty, we propose to use $e_{i}^{\tt agg*}$ as defined below for evidence aggregation.
$e_{i}^{\tt agg*}=\frac{1}{n_{i}}\sum_{k=1}^{n_{i}}\binom{n_{i}}{k}^{-1}\sum_{\mathcal{S}_{ki}\in\mathcal{B}_{ki}}e_{i,\mathcal{S}_{k,i}},$ (5)
where $\mathcal{B}_{ki}$ is the set of all $k$ element subsets of $\mathcal{N}_{i}$. The idea is to try all possible $\pi^{k}A_{i,\mathcal{S}_{ki}}$ and then take average. Since $\pi^{k}A_{i,\mathcal{S}_{ki}}$ are $e$-values, their average is also an $e$-value. The form of $e_{i}^{\tt agg*}$ in Equation (5) also appears in Vovk and Wang (2024) as “U-statistics” $e$-values. We denote the procedure that applies the e-BH procedure on $(e_{1}^{\tt agg*},\ldots,e_{m}^{\tt agg*})$ as IRT. In the real data analysis and numerical experiments of sections G and F we take $\pi_{j}=0.5$ for all $j\in[d]$, as is often assumed in the literature (Jin and Cai, 2007).
Remark 1.
Note that from Theorem 3, IRT guarantees valid FDR control under (1) exchangeability of the study-specific summary statistics, (2) symmetry of the study-specific decision rule and (3) independence of the summary statistics for each testing problem. If the data underlying the null hypotheses are independent, then the resulting $p$-values, which are common summary statistics, can be exchangeable, regardless of effect heterogeneity. A key aspect here is that under the null, such $p$-values should follow a Uniform(0,1) distribution. In contrast, if the data under the null hypotheses exhibit, for instance, spatial dependence then the corresponding $p$-values may not be exchangeable. In Section F.3 we assess the numerical performances of IRT when the assumptions underlying Lemma 1 and Theorem 3 are violated. In particular, we find that IRT is relatively robust to the exchangeability assumption on the summary statistics in Lemma 1.
Remark 2.
The scaling by $\pi_{j}$ in Lemma 1 is a theoretical requirement for constructing a valid e-value, but it also provides a crucial and desirable statistical calibration. The magnitude of evidence from a discovery should be calibrated by how difficult it was to make. Most modern multiple testing procedures are inherently data-adaptive, and the bar for declaring a finding significant is often lower in a dense-signal environment. In the BH procedure, for instance, a higher proportion of true signals makes the procedure more powerful, leading to a less stringent p-value threshold that satisfies the FDR criterion. This implies that a hypothesis rejected in a dense-signal study may not have needed to achieve as small a p-value as one rejected in a sparse-signal study. It is therefore natural and statistically sound for discoveries from dense-signal settings to contribute a lower evidence weight. The scaling by $\pi_{j}$ in our IRT framework is the mechanism that performs this automatic and desirable calibration.
We now return to the example considered in Section A but instead use IRT to pool the inferences from the two studies. The green and the gray shaded regions in the left panel of Figure 4 depicts the overall FDR level $\alpha$ required for the e-BH procedure to reject $H_{0i}$ as a function of $\alpha_{1}$ and with $\alpha_{2}=0.05$. Clearly, IRT is able to reject $H_{0i}$ even when $\alpha<\min(\alpha_{1},\alpha_{2})$. This represents a stark contrast to IRT which requires $\alpha$ to stay in the gray region to reject $H_{0i}$. The reason for this distinction is related to the fact that, courtesy Theorem 3, the aggregation scheme underlying IRT is more powerful than the arithmetic mean. Note that the calculations for the left panel directly follow from the IRT procedure which, in this example, rejects $H_{0i}$ at FDR level $\alpha$ if $\alpha\geq 8\alpha_{1}\alpha_{2}/(\alpha_{1}+\alpha_{2}+m)$. While this example relies on a relatively simple setting, our numerical experiments in Section F of the supplement confirm the broader conclusion that, in general, IRT is not powerful when $\alpha<\min_{j\in[d]}\alpha_{j}$ while, under the conditions of Theorem 3, IRT is powerful in this setting.
Remark 3.
Motivated by the setting in the left panels of figures 3 and 4, suppose there are $d=4$ studies, each testing the same set of $m=5$ null hypotheses. The first two studies are testing at level $\alpha_{1}$ while the remaining two at $\alpha_{2}$. Consider a setting where the first two studies are using exactly the same data and both reject only the $i$-th null hypothesis at FDR level $\alpha_{1}$. Other two studies are independent such that Theorem 3 holds, but both continue to reject the same $i$-th null hypothesis at FDR level $\alpha_{2}$. So $\delta_{ij}=1$ and $\|\bm{\delta}_{j}\|_{0}=1$ for all $j\in[d]$. To pool inferences across these $d$ studies, IRT and IRT can be used in conjunction. Specifically, the aggregated evidence indices from the first two studies are constructed using IRT, denoted $e_{i}^{\tt agg}$ (Equation (4)) and IRT is employed to aggregate the evidence indices from the remaining two studies, denoted $e_{i}^{\tt agg*}$ (Equation (5)). Denote $e_{i}^{\tt agg,H}=(1/2)(e_{i}^{\tt agg}+e_{i}^{\tt agg*})$ as the hybrid aggregated evidence indices. Note that $\{e_{i}^{\tt agg,H}\}_{i=1}^{m}$ are a set of compound $e$-values under $\mathcal{H}_{0}$. When $\{e_{i}^{\tt agg,H}\}_{i=1}^{m}$ are used as inputs to the e-BH procedure, the gray and the green shaded regions in the right panel of Figure 4 depict the overall FDR level $\alpha$ required for e-BH to reject $H_{0i}$ as a function of $\alpha_{1}$ and with $\alpha_{2}=0.05$. Clearly, this hybrid scheme is able to reject $H_{0i}$ even when $\alpha<\min(\alpha_{1},\alpha_{2})$. In contrast, the gray shaded region reveals that IRT has no power unless $\alpha>\min(\alpha_{1},\alpha_{2})$.
The above example represents a practical setting where often additional information regarding data-sharing or the use of auxiliary side information for multiple testing is available for the $d$ studies. For instance, suppose prior knowledge dictates that a set of $d_{1}\subset[d]$ studies share data amongst themselves, while the conditions of Theorem 3 hold for the remaining set of $d_{2}=[d]\setminus d_{1}$ studies. Denote the aggregated evidence indices across the $d_{1}$ studies derived from IRT as $\{e_{i}^{\tt agg,d_{1}}\}_{i\in\mathcal{M}}$ and those derived from IRT across the $d_{2}$ studies as $\{e_{i}^{\tt agg*,d_{2}}\}_{i\in\mathcal{M}}$. Then, the hybrid aggregated evidence indices $e_{i}^{\tt agg,H}=(1/d)(|d_{1}|e_{i}^{\tt agg,d_{1}}+|d_{2}|e_{i}^{\tt agg*,d_{2}})$ are a set of compound $e$-values under $\mathcal{H}_{0}$ and the e-BH procedure guarantees FDR control at level $\alpha$ when $\{e_{i}^{\tt agg,H}\}_{i=1}^{m}$ are used as inputs. We call this procedure IRT H and evaluate its empirical performance in Section F of the supplement.
Appendix C Proofs
C.1 Proof of Theorem 1
Proof.
Based on the evidence construction in Equation (3), we have
$\sum_{i\in\mathcal{H}_{0j}}\mathbb{E}(e_{ij})$ $=\dfrac{m_{j}}{\alpha_{j}}~\mathbb{E}\left[\dfrac{\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}}{\|\bm{\delta}_{j}\|_{0}\vee 1}\right]=\dfrac{m_{j}}{\alpha_{j}}~\text{FDR}(\bm{\delta}_{j})\leq m_{j},$
C.2 Proof of Theorem 2
Proof.
We have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}^{\texttt{agg}}]$ $=\dfrac{1}{d}\sum_{i\in\mathcal{H}_{0}}\sum_{j=1}^{d}\Bigl{\{}\dfrac{m_{j}}{\alpha_{j}}~\mathbb{E}\left[\dfrac{\delta_{ij}}{\max(\|\bm{\delta}_{j}\|_{0},1)}\right]~\mathbb{I}(i\in\mathcal{M}_{j})+\mathbb{I}(i\notin\mathcal{M}_{j})\Bigr{\}}$
$=\dfrac{1}{d}\sum_{j=1}^{d}\Bigl{\{}\dfrac{m_{j}}{\alpha_{j}}~{\rm FDR}(\bm{\delta}_{j})+\sum_{i\in\mathcal{H}_{0}}\mathbb{I}(i\notin\mathcal{M}_{j})\Bigr{\}}$
$\leq\dfrac{1}{d}\sum_{j=1}^{d}\Bigl{\{}m_{j}+m-m_{j}\Bigr{\}}=m,$
which completes the proof. ∎
C.3 Proof of Lemma 1
Proof.
@@ C.3 Proof of Lemma 1
C.4 Proof of Theorem 3
Proof.
Note that for $i\in\mathcal{H}_{0}$, $\mathbb{E}_{H_{0i}}(e_{i,\mathcal{S}_{ki}})=\prod_{j\in\mathcal{S}_{ki}}\pi_{j}\mathbb{E}_{H_{0i}}(e_{ij})\leq\prod_{j\in\mathcal{S}_{ki}}\pi_{j}\dfrac{m_{j}}{|\mathcal{H}_{0j}|}\leq 1$. ∎
Appendix D Additional technical details
D.1 Connections to existing aggregation and derandomization procedures